# Bayesian Model Averaging in Causal Instrumental Variable Models

### v4

@@ 8 Conclusion
age 0.704 0.809 0.781 0.903 0.999 1.0 0.801
agesq 0.298 0.607 0.233 0.774 0.367 1.0 0.329
nearc2 0.12 0.004 0.397 0.018 0.488 0.15 0.823
nearc4 0.001 0.052 0.013 0.102 0.02 0.457 0.083
momdad14 0.003 0.023 0.016 0.041 0.049 0.317 0.094
sinmom14 0.003 0.004 0.007 0.01 0.11 0.355 0.079
step14 0.004 0.291 0.032 0.532 0.076 0.912 0.195
black 0.999 0.011 1.0 0.039 1.0 0.292 1.0
south 1.0 0.006 1.0 0.018 1.0 0.108 1.0
smsa 1.0 0.944 1.0 0.958 1.0 0.982 1.0
married 1.0 0.879 1.0 0.983 1.0 0.993 1.0
reg662 0.001 0.002 0.007 0.011 0.025 0.135 0.106
reg663 0.042 0.007 0.163 0.017 0.25 0.126 0.555
reg664 0.007 0.016 0.04 0.044 0.144 0.464 0.169
reg665 0.003 0.022 0.008 0.032 0.029 0.257 0.094
reg666 0.0 0.004 0.011 0.015 0.032 0.16 0.08
reg667 0.001 0.009 0.005 0.026 0.03 0.269 0.094
reg668 0.244 0.005 0.611 0.009 0.841 0.347 0.771
reg669 0.001 0.006 0.016 0.019 0.036 0.21 0.105
fatheduc 0.006 1.0 0.14 1.0 0.073 1.0 0.105
motheduc 0.008 1.0 0.056 1.0 0.085 1.0 0.796
Table S.7: Returns to schooling: Posterior inclusion probabilities for the Card, (1995) dataset including parental education, based on the reduced sample of $n=2,215$ observations.
Figure S.2: Returns to schooling: Posterior distributions of the treatment effect of education (Rao-Blackwellized for gIVBMA; Kernel density estimate for IVBMA) and the covariance ratio $\sigma_{yx}/\sigma_{xx}$ based on the Card, (1995) dataset without parental education but only using the observations that have complete entries for parental education ($n=2,215$). The algorithm was run for $50,000$ iterations, the first $5,000$ of which were discarded as burn-in. IVBMA includes a point mass at zero of $22.8\%$.
gIVBMA (hyper-$g/n$) gIVBMA (BRIC) IVBMA BMA (hyper-$g/n$)
L M L M L M L
age 0.99 0.229 0.501 0.739 1.0 1.0 0.805
agesq 0.001 0.157 0.521 0.662 0.22 1.0 0.306
nearc2 0.135 0.142 0.59 0.131 0.501 0.37 0.874
nearc4 0.001 0.024 0.012 0.073 0.032 0.4 0.082
momdad14 0.001 0.014 0.01 0.033 0.07 0.277 0.09
sinmom14 0.001 0.005 0.008 0.019 0.1 0.321 0.103
step14 0.002 0.106 0.027 0.209 0.09 0.658 0.186
black 0.006 1.0 0.055 1.0 0.384 1.0 1.0
south 0.01 0.996 0.282 0.82 0.402 0.797 1.0
smsa 0.002 1.0 0.074 1.0 0.289 1.0 1.0
married 1.0 0.999 1.0 1.0 1.0 1.0 1.0
reg662 0.001 0.007 0.011 0.009 0.041 0.135 0.113
reg663 0.031 0.017 0.347 0.04 0.418 0.195 0.57
reg664 0.001 0.003 0.024 0.016 0.11 0.351 0.155
reg665 0.001 0.001 0.01 0.018 0.03 0.193 0.096
reg666 0.001 0.004 0.007 0.021 0.033 0.143 0.081
reg667 0.001 0.003 0.011 0.011 0.031 0.13 0.066
reg668 0.057 0.011 0.3 0.028 0.747 0.614 0.742
reg669 0.001 0.003 0.012 0.019 0.039 0.165 0.097
Table S.8: Returns to schooling: Posterior inclusion probabilities for the Card, (1995) dataset without parental education but only using the observations that have complete entries for parental education ($n=2,215$).
Figure S.3: Returns to schooling: Posterior distribution of the treatment residual variance $\sigma_{xx}$ for gIVBMA (BRIC and hyper-$g/n$) and IVBMA.
@@
Method (a) (b) (c)
gIVBMA (hyper-$g/n$) 0.432 0.442 0.441
gIVBMA (BRIC) 0.432 0.439 0.442
BMA (hyper-$g/n$) 0.431 0.439 0.44
IVBMA 0.432 0.441 0.441
TSLS 0.698 0.591 0.72
Table S.9: Returns to schooling: The mean LPS calculated over each fold of the Card, (1995) data (a) without parental education ($n=3,003$), (b) with parental education ($n=2,215$), and (c) without parental education only using the shared set of observations ($n=2,215$) in a 5-fold cross-validation procedure.
S.5.2 Additional Example: Maternal smoking and birthweight
As an additional application, we revisit the effect of maternal smoking during pregnancy on birth weight using data from Mullahy, (1997). The smoking habits of mothers are likely to be correlated with other unobserved determinants of the child’s birth weight, potentially leading to endogeneity. The dependent variable is birth weight in ounces, and the endogenous regressor is the typical number of cigarettes smoked daily during the pregnancy. As instruments, we use the father’s years of education, the mother’s years of education, the cigarette price in their home state, and the excise tax on cigarettes in their home state. These have a plausible effect on the mother’s smoking habits but are unlikely to affect the child’s birth weight directly. Further explanatory variables are birth order, the child’s sex, and race. We also include all pairwise interactions between the instruments and exogenous covariates as potential covariates (they could end up being selected as instruments). In total, this yields $19$ potential instruments and covariates. After removing missing values, the dataset contains $1,191$ observations.
The data records both birth weight and the number of cigarettes smoked daily as integers. However, the birth weight numbers are reasonably large, such that a Gaussian approximation could work well. This gives us two modelling options: a Poisson with a latent log-normal parameter or a Gaussian log-linear model. The latter uses the natural logarithm of birth weight as an outcome to ensure the interpretation is comparable across these two models. We fit both models with hyper-$g/n$ priors and compare the results. In both cases, the number of cigarettes (the only endogenous variable) is modeled as a Poisson with a latent log-normal parameter. We also compare this to naive BMA (also under a hyper-$g/n$ prior) and IVBMA, both applied to the log-linear Gaussian model. Note that IVBMA also treats the number of cigarettes in the treatment model as Gaussian.
Figure S.4: Maternal smoking and birthweight: The marginal posterior distribution (Rao-Blackwellised) of the effect of smoking an additional cigarette a day on birth weight. The algorithm was run for 5,000 iterations (the first 1,000 of which were discarded as burn-in; 50,000 and 10,000 for IVBMA). IVBMA has point mass at zero (99.299.2% of the posterior mass) indicated by the vertical line. The continuous part is not rescaled for better visibility.
Figure S.4 shows the marginal posterior for the cigarettes coefficient in the Poisson and the Gaussian model. They are both centered at just above 0.005-0.005, but the Poisson is slightly tighter. For each additional cigarette smoked per day, we would expect a decrease in birth weight of around $0.5\%$. IVBMA does not select the endogenous variable most of the time and, therefore, has most of its posterior mass on zero. Both the Gaussian and the Poisson models always select mothers’ years of education as an instrument, while other instruments are only sporadically included. IVBMA, on the other hand, selects the father’s years of education in almost all treatment models, while the mother’s education is only included about two-thirds of the time.
Again, we compare the predictive performance on this dataset using the LPS in a cross-validation procedure. For the log-linear models, we adjust the posterior predictive by the Jacobian of $y\mapsto\log(y)$ to be on the same scale as the Poisson log-normal model. So, the approximate LPS for the log-linear models is
LPS $\approx-\frac{1}{n^{*}}\sum_{i=1}^{n^{*}}\log\left(\frac{1}{S}\sum_{s=1}^{S}\frac{p(\log y_{i}^{*}\mid X_{i}^{*},\theta^{s})}{|y_{i}^{*}|}\right)$
$=-\frac{1}{n^{*}}\sum_{i=1}^{n^{*}}\log\left(\frac{1}{S}\sum_{s=1}^{S}p(\log y_{i}^{*}\mid X_{i}^{*},\theta^{s})\right)+\frac{1}{n^{*}}\sum_{i=1}^{n^{*}}\log|y_{i}^{*}|.$
Table S.10 presents the results of the cross-validation exercise. The Poisson log-normal model achieves the lowest mean LPS.
Method Mean LPS
BMA 4.497
gIVBMA (PLN) 4.475
gIVBMA (Gaussian) 4.504
IVBMA (KL) 4.521
TSLS 4.522
Table S.10: Maternal smoking and birthweight: The mean LPS calculated over each fold of the birthweight data in a 5-fold cross-validation procedure.

### v5

@@ 8 Conclusion
exper 1.0 1.0 1.0 1.0 1.0 1.0 1.0
expersq 1.0 0.001 1.0 0.001 0.389 0.006 1.0
nearc2 0.128 0.032 0.111 0.0 0.324 0.106 0.754
nearc4 0.0 0.12 0.001 0.096 0.022 0.42 0.09
momdad14 0.0 0.007 0.0 0.0 1.0 0.144 0.154
sinmom14 0.004 0.003 0.0 0.007 0.799 0.258 0.166
step14 0.01 1.0 0.0 1.0 1.0 1.0 0.284
black 1.0 0.2 1.0 0.21 1.0 0.682 1.0
south 1.0 0.018 1.0 0.052 1.0 0.094 1.0
smsa 1.0 0.784 1.0 0.805 1.0 0.972 1.0
married 1.0 0.421 1.0 0.406 1.0 0.728 1.0
reg662 0.0 0.026 0.0 0.0 0.035 0.123 0.12
reg663 0.066 0.008 0.046 0.004 0.126 0.084 0.605
reg664 0.011 0.001 0.009 0.016 0.276 0.273 0.314
reg665 0.0 0.0 0.0 0.0 0.026 0.118 0.132
reg666 0.026 0.012 0.0 0.0 0.028 0.25 0.115
reg667 0.0 0.0 0.0 0.014 0.037 0.144 0.148
reg668 0.364 0.016 0.342 0.03 0.9 0.449 0.846
reg669 0.0 0.137 0.0 0.175 0.026 0.813 0.097
fatheduc 0.015 1.0 0.018 1.0 0.01 1.0 0.095
motheduc 0.012 1.0 0.004 1.0 0.024 1.0 0.826
Table S.7: Returns to schooling: Posterior inclusion probabilities for the Card, (1995) dataset only using observations that have complete parental education values ($n=2,215$).
Figure S.6: Returns to schooling: The implied posterior distribution of the number of valid and relevant instruments for gIVBMA with the inverse Wishart prior (IW), the Cholesky ($\omega_{a}=0.1$) prior, and IVBMA.
@@
Method Imputed ($n=3,003$) Complete ($n=2,215$)
gIVBMA (IW) 0.425 0.434
gIVBMA ($\omega_{a}=0.1$) 0.424 0.434
BMA 0.426 0.435
IVBMA 0.439 0.442
TSLS 0.584 0.584
Table S.8: Returns to schooling: The mean LPS calculated over each fold of the imputed data ($n=3,003$) and the complete parental education data ($n=2,215$) in a 5-fold cross-validation procedure.